OpenStudy (anonymous):

Computer Craft pays its sales associates a 15% commission on every sale. Barry sold a laptop computer for $1,900 and a printer for $300. How much money did Barry earn on the sale of the items?

OpenStudy (jazzyfa30):

ok first add up the amounts what do you get??

OpenStudy (anonymous):

wait...on every sale....isn't that two sales ?

OpenStudy (jazzyfa30):

yes so if they add up both amounts then multiply by .30 that will there answer

OpenStudy (mathstudent55):

Multiply the sum by 0.15

OpenStudy (jazzyfa30):

no @mathstudent55 the Q states each sale and there are 2 sales so u multiply by .30

OpenStudy (anonymous):

wouldn't you multiply the sum by .30

OpenStudy (jazzyfa30):

yes

OpenStudy (anonymous):

were these items purchased separately ?

OpenStudy (mathstudent55):

The commission is 15%. 15% of 1900 + 15% of 300 = 0.15 * 1900 + 0.15 * 300 Factor out 0.15: 0.15 * ( 1900 + 300) 15% * (1900 + 300) In other words, take 15% of the sum.

OpenStudy (jazzyfa30):

thats still .30 either way it is the same answer

OpenStudy (jazzyfa30):

and there is no need to factor

OpenStudy (mathstudent55):

If you multiply the sum by 0.30, that means you are making the commission 30% on each item. You can't add the commission percentages. You can only add the prices of the items.

OpenStudy (anonymous):

I would just do the .30 with the sum

OpenStudy (mathstudent55):

@jazzyfa30 No it's still not .30 and it never was.

OpenStudy (jazzyfa30):

thats the same thing i think @kelliegirl33

OpenStudy (anonymous):

so it is just 15% of EACH sale and not 30% of the total ?

OpenStudy (jazzyfa30):

k so if we do one item at a time then add them together were still putting it in as .30o

OpenStudy (mathstudent55):

@jazzyfa30 & @kelliegirl33 Stop writing for a minute and follow along what I am writing below.

OpenStudy (mathstudent55):

Let me ask you this way. Imagine the employee makes a 50% commission. He sells 2 items. Both items cost $100 each. According to what you are saying, the commission now should be 100% of $200, right?

OpenStudy (anonymous):

it would be half... 50% of 100 + 50% of 100

OpenStudy (jazzyfa30):

which would make it all half price

OpenStudy (anonymous):

cost would be 200 and commission would be 100...right ?

OpenStudy (mathstudent55):

That is correct. But abovem regarding the 15% commission, you said that it would be 30% of the sum which in this case would correspond to 100% of $200.

OpenStudy (mathstudent55):

That is correct. But above regarding the 15% commission, you said that it would be 30% of the sum which in this case would correspond to 100% of $200.

OpenStudy (anonymous):

what about 15% of 1900 and 15% of 300...but if you didn't factor would you get the same answer as you would by factoring ?

OpenStudy (jazzyfa30):

ok i get that but if we did each item 1 at a time we would still technically be puttig the answer as 30% of the overall amount do u see what im getting at??

OpenStudy (mathstudent55):

In my example (50% commission) isn't 50% of 100 + 50% of 100 the same as 50% of (100 + 100) = 50% of 200?

OpenStudy (anonymous):

oh...yes it is

OpenStudy (jazzyfa30):

but the answer either way would be 100

OpenStudy (mathstudent55):

It';s is still not 30% of anything 15% of 1900 + 15% of 300 = 15% of (1900 + 300) There is not 30% anywhere.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

so don't add the percents, just the sums

OpenStudy (jazzyfa30):

ok but the answer will stilll be the same either way just like in your example

OpenStudy (anonymous):

I got it now....

OpenStudy (mathstudent55):

Correct. From my 50% commission example, you are not adding the comission percentages and saying it's 100% total commssison. Why do you keep insisting that the 15% commssions add together to make a 30% commssion?

OpenStudy (anonymous):

30% of 2200 is not the same as 15% of 2200

OpenStudy (anonymous):

I understand what you are saying now mathstudent....it makes sense to me now

OpenStudy (jazzyfa30):

yeah i understand too and @Babez_1999 ur answer is 330

OpenStudy (mathstudent55):

In my example, if you added the percents of the commissions, then the commission would be 100%. 100% of the $200 in sales ($100 + $100) would be $200 in commission, but you know the commission is only $100. So you know that the percentages do not add together.

OpenStudy (anonymous):

yea...that makes sense

OpenStudy (mathstudent55):

@jazzyfa30 I know the problem, does not require factoring, but those steps above that I showed in which I factored were to prove mathematically that if you take 15% commissions of many items and add them together, it's the same as taking 15% commission of the sum.

OpenStudy (jazzyfa30):

After avery longgggggggggggggggg debate mathstudent proved me wrong and i actually appriciated it thanks @mathstudent55 u taught me sumn new.

OpenStudy (anonymous):

me too :)

OpenStudy (jazzyfa30):

yea i see that

OpenStudy (mathstudent55):

15% of 1900 = 0.15 * 1900 = 285 15% of 300 = 0.15 * 300 = 45 ------(add) 330 15% of (1900 + 300) = 15% of 2200 = 330

OpenStudy (mathstudent55):

Great

OpenStudy (anonymous):

thanks mathstudent

OpenStudy (jazzyfa30):

thanks

OpenStudy (mathstudent55):

wlcm

OpenStudy (mathstudent55):

BTW, with this, you can save a lot of time. Let's say you go to a store and every item is on sale, 10% off. You are buying 20 items and want to find the total price. Instead of taking off 10% of each item, then adding all the dicounted prices together, you can simply just add all original prices together, then take 10% off from the sum. You only calculate a percentage once.